### v4

# Conformal Temporal Logic Planning using Large Language Models

Abstract
This paper addresses planning problems for mobile robots. We consider missions that require accomplishing multiple high-level sub-tasks, expressed in natural language (NL), in a temporal and logical order. To formally define the mission, we treat these sub-tasks as atomic predicates in a Linear Temporal Logic (LTL) formula. We refer to this task specification framework as LTL-NL. Our goal is to design plans, defined as sequences of robot actions, accomplishing LTL-NL tasks. This action planning problem cannot be solved directly by existing LTL planners because of the NL nature of atomic predicates. To address it, we propose HERACLEs, a hierarchical neuro-symbolic planner that relies on a novel integration of (i) existing symbolic planners generating high-level task plans determining the order at which the NL sub-tasks should be accomplished; (ii) pre-trained Large Language Models (LLMs) to design sequences of robot actions based on these task plans; and (iii) conformal prediction acting as a formal interface between (i) and (ii) and managing uncertainties due to LLM imperfections. We show, both theoretically and empirically, that HERACLEs can achieve user-defined mission success rates. Finally, we provide comparative experiments demonstrating that HERACLEs outperforms LLM-based planners that require the mission to be defined solely using NL. Additionally, we present examples demonstrating that our approach enhances user-friendliness compared to conventional symbolic approaches.
Index Terms:
Neuro-symbolic Planning, Linear Temporal Logic, Large Language Models, Conformal Prediction.
I Introduction
Designing autonomous agents with task planning capabilities is a long standing goal in robotics and automation [1, 2, 3, 4]. Achieving this goal requires the development of mission specification frameworks and planning algorithms capable of generating plans—sequences of robot actions—to accomplish assigned missions. These frameworks should allow practitioners to define missions unambiguously and in a user-friendly manner, while planners need to exhibit computational efficiency and be supported by correctness guarantees.
Recently, several task planners have been proposed that can design correct-by-construction plans for complex missions with specified temporal and logical requirements using Linear Temporal Logic (LTL) [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. However, this framework demands a significant amount of expertise and manual effort to rigorously define complex tasks. This is because defining LTL tasks requires specifying multiple atomic predicates (i.e., Boolean variables) to model desired low-level robot configurations and coupling them using temporal/Boolean operators. Additionally, the more complex the task requirements, the larger the number of predicates and temporal/logical operators needed to define the corresponding LTL formula. This not only compromises the user-friendliness of the framework but also increases the computational cost of designing robot plans [14].
On the other hand, Natural Language (NL) has emerged as a more user-friendly method for specifying robot missions. Early research in NL-based planning primarily focused on mapping NL to planning primitives [30, 31, 32, 33, 34] using statistical machine translation [35] to identify data-driven patterns for translating free-form commands into a formal language defined by a grammar. However, these approaches were limited to structured state spaces and simple NL commands. Motivated by the remarkable generalization abilities of pre-trained Large Language Models (LLMs) across diverse task domains [36, 37, 38, 39], there has been increasing attention on utilizing LLMs for NL-based planning [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. While LLM-based planning methods are more user-friendly than LTL-based planners, they have two main limitations: (i) they lack mission performance guarantees; and (ii) their ability to design correct plans deteriorates with increasing task complexity [57].
In this paper, we propose an alternative approach to mission specification and robot planning aimed at addressing some of the limitations observed in current LTL and NL planning frameworks. Our approach involves tightly coupling LTL and NL to leverage their respective strengths. First, we discuss our proposed mission specification framework. We model a complex task as comprised of a collection of sub-tasks, with temporal and logical relationships among them represented using an LTL specification. The sub-tasks thus serve as atomic predicates in an LTL formula, with the associated predicate being true if the sub-task is successfully accomplished, and false otherwise. Our key departure from conventional LTL specifications is that we specify each such sub-task in NL, rather than directly grounding it in the low-level configuration space; see e.g., [10, 11, 14, 15, 24]. For example, consider a simple task in which the robot first needs to deliver a bottle of water to the kitchen table, and only afterwards deliver a coke to the office desk. We can define two NL sub-tasks, ‘deliver a bottle of water to the kitchen table’ and ‘deliver a coke to the office desk’, and use an LTL formula to express the logical and temporal relationship between them. The advantage of this framework over conventional LTL is that LTL-NL formulas require fewer predicates and temporal/logical operators while capturing the same task, making them more user-friendly to define. In comparison to NL instructions (which are more user-friendly to define than LTL-NL commands), our framework enables the automatic decomposition of the overall mission (i.e., an LTL-NL specification) into multiple NL-based sub-tasks using existing symbolic planners [19, 20, 26]. It is important to note that such a decomposition is highly challenging to perform in a correct-by-construction fashion for missions expressed exclusively using NL. As we will discuss later, task decomposition plays a critical role in enabling our planning framework to consistently outperform existing LLM-based planners, especially as mission complexity increases.
Figure 1: This paper proposes HERACLEs, a neuro-symbolic planning framework for mobile robots with LTL-NL missions. The framework starts with a symbolic task planner that generates the next language-based sub-task that should be accomplished to make mission progress. Pre-trained LLMs are then deployed to design robot plans for that sub-task. The LLM planner seeks assistance from the symbolic planner either when the robot plan cannot be executed due to environmental constraints or when it is uncertain about how to accomplish that sub-task. Uncertainty of the LLM is quantified using conformal prediction (CP). If the LTL task planner cannot assist, then help from users is requested (not shown). The mission in this figure requires the robot to eventually deliver a pen to location D but only after either an apple or a drink has been delivered to D. The symbolic planner first generates the sub-task ‘deliver a drink to E’. Since the prediction set using CP is non-singleton (including the actions ‘Go to location A’ and ‘Go to location D’), the LLM-based planner asks the symbolic planner to generate an alternative sub-task to make progress. The new task is ‘deliver an apple to E’; see also Example II.4.
Second, we address the challenge of designing robot plans that ensure correctness with respect to LTL-NL formulas. Specifically, we consider a mobile robot equipped with various skills (e.g., mobility, manipulation, and sensing) tasked with missions expressed as LTL-NL specifications. Each predicate in the LTL-NL formula is satisfied if an NL-based sub-task is accomplished, requiring the robot to apply its skills to various semantic objects and regions of interest in the environment. For instance, the sub-task ‘deliver a bottle of water to the kitchen table’ requires the robot to move to the location where the bottle is, grasp it, move to the kitchen table, and release the object. Observe that a predicate in an LTL-NL formula is satisfied by a sequence of robot configurations/actions. While we assume that the locations and semantic labels of these objects are known, the geometric environmental structure is initially unknown. The latter may hinder the completion of certain NL-based sub-tasks (e.g., when an object is inaccessible). Our goal is to design robot plans, defined as a sequence of robot actions, satisfying the assigned LTL-NL missions.
To address this task planning problem, we propose a hierarchical neuro-symbolic planner that harnesses the complementary strengths of LLMs and symbolic planners. Specifically, our proposed planner leverages existing LTL planners that dynamically determine, based on the current mission status, the next NL-based sub-task the robot should accomplish to make progress in the mission. In our implementation, we select the symbolic planner proposed in [19, 20] due to its computational efficiency; however, we emphasize that any other symbolic planner can be used. Then, we employ pre-trained LLMs to generate robot plans satisfying the NL sub-task announced by the symbolic planner [37, 58]. If the geometric environmental structure prevents the robot from executing these plans (e.g., obstacles blocking the entrance to regions of interest), the LTL task planner will generate alternative sub-tasks (if they exist) as in [19, 20]. A challenge here is that LLMs tend to hallucinate, i.e., to confidently generate incorrect outputs [59]. Such false confidence in incorrect outputs poses a significant challenge in LLM-based robot planning. To reason about the correctness of the LLM-generated plans, inspired by [41, 60], we leverage conformal prediction (CP) [61, 62, 63], a statistical tool for uncertainty quantification in black-box models that has been recently applied to various safe autonomy tasks [64, 65, 60, 41, 66, 67, 68, 69, 70]. CP constructs on-the-fly prediction sets that contain the correct robot action with user-specified confidence. This allows the LLM to determine when it is uncertain about its predictions. In cases of high uncertainty, indicated by non-singleton prediction sets, the robot seeks assistance as follows. First, a help-message is sent to the symbolic planner requesting alternative sub-tasks to make mission progress. If such a sub-task exists, it is forwarded to the LLM planner. Otherwise, the robot asks for help from a user about how to proceed. This formal interface between the symbolic planner and the pre-trained LLM gives rise to our planner, called HERACLEs, for HiERArchical ConformaL natural languagE planner; see Fig. 1. The generated plans are executed using existing low-level controllers. We show both theoretically and empirically that HERACLEs can achieve user-specified theoretical mission success rates while requiring low help rates.
Related works: (i) As discussed earlier, a substantial body of symbolic planners exists that can design plans satisfying LTL tasks [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. We emphasize that these planners cannot solve the considered LTL-NL planning problem due to the NL nature of the predicates. Specifically, these works can design a sequence of NL-based predicates that need to be satisfied to accomplish the LTL-NL mission; however, they cannot design a sequence of robot actions (e.g., ‘grab the bottle of water at location A’, ‘go to the kitchen table’, ‘drop off the bottle’) that satisfies a NL-based predicate (e.g., ‘deliver a bottle of water to the kitchen table’), let alone an LTL-NL formula, as they cannot process NL. We note again that our approach is tightly integrated with these symbolic planners as they are used to generate a sequence of NL sub-tasks that should be accomplished to satisfy an LTL-NL formula. (ii) Related are also recent neural planners that can design plans for tasks, described solely in NL, using pre-trained LLMs; see e.g., [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]; a recent survey can be found in [71, 72, 73]. However, the ability of LLMs to design correct plans drops significantly as the mission complexity increases as shown in [57, 74] as well as in our experiments. Additionally, as discussed earlier, the majority of these planners lack correctness and mission completion guarantees. Among these works, probably the closest one is [41] as it also applies CP for uncertainty alignment of LLMs allowing robots to ask for help from users in order to achieve desired mission completion rates. However, as shown in our comparative experiments, HERACLEs requires significantly lower help rates than [41] due to its task decomposition feature enabled by the LTL-NL specification. (iii) Neuro-symbolic planners that integrate LLMs with symbolic planners have also been proposed recently in [57, 75, 76, 77, 78, 79, 80]. Their key idea is to employ LLMs to translate NL commands into temporal logic specifications. Once this translation is completed, existing symbolic planners can be employed to design correct plans. A key limitation is that the translation process is not supported by correctness guarantees. As a result, the resulting plan may not satisfy the original NL instructions. In contrast, HERACLEs employs LLMs and symbolic planners in a fundamentally different way while CP allows the design of plans to achieve user-specified mission completion rates.
Comparative Evaluations: We provide extensive comparative experiments showing that HERACLEs outperforms state-of-the-art LLM-based planners in terms of planning performance. In these experiments, we manually translate LTL-NL missions into NL instructions which serve as inputs to these planners. First, we compare HERACLEs against the LLM-based planners proposed in [40, 50, 81], which, however, do not utilize CP and do not allow robots to ask for help. Our comparative experiments show that the ability of these approaches to design correct plans tends to decrease as the number of temporal and logical requirements in the mission increases. To the contrary, the planning performance of HERACLEs consistently outperforms these baselines (even when the help mode is deactivated). Second, we compare our planner against a conformalized planner [41] that also allows robots to ask for help to achieve desired mission success rates. Our empirical analysis demonstrates that HERACLEs achieves significantly lower help rates than [41] to attain the same mission success rate. We attribute the performance gap between these baselines [40, 50, 81, 41] and HERACLEs primarily to the LTL-NL framework and the symbolic planner decomposing missions into multiple smaller/simpler NL-based sub-tasks that LLMs can handle more efficiently. Finally, we provide examples demonstrating the user-friendliness of LTL-NL formulas over the corresponding LTL specifications. Specifically, we show that LTL-NL formulas require significantly fewer predicates than an LTL formula modeling the same task. This difference in the ‘length’ of LTL and LTL-NL formulas becomes more pronounced as the complexity of the mission requirements increases.
Contribution: The contribution of the paper can be summarized as follows. First, we propose a new task specification approach, called LTL-NL, to define complex high-level robot tasks. The advantage of LTL-NL over conventional LTL is its user-friendliness as it requires fewer predicates to define the same task while the predicates can be defined using NL, rather than low-level system configurations. Second, we introduce HERACLEs, a new neuro-symbolic planner that integrates existing symbolic planners with pre-trained LLMs to design plans satisfying missions encoded as LTL-NL formulas. Third, we show, both theoretically and empirically, that HERACLEs can achieve user-specified mission success rates through a formal interface, employing CP, between the LLM and the symbolic planner. Fourth, we provide extensive comparative experiments highlighting that HERACLEs outperforms state-of-the-art LLM-based planners in terms of its ability to design correct plans and achieve lower help rates.
II Problem Formulation
Robot System and Skills: Consider a robot governed by the following deterministic and known dynamics:
${\mathbf{p}}(t+1)={\mathbf{f}}({\mathbf{p}}(t),{\mathbf{u}}(t)),$ (1)
where ${\mathbf{p}}(t)$ stands for the state (e.g., position and orientation) of the robot, and ${\mathbf{u}}(t)$ stands for control input at discrete time $t$. We assume that the robot state ${\mathbf{p}}(t)$ is known for all time instants $t\geq 0$. The robot has $A>0$ number of abilities/skills collected in a set ${\mathcal{A}}\in\{1,\dots,A\}$. Each skill $a\in{\mathcal{A}}$ is represented as text such as ‘take a picture’, ‘grab’, or ‘move to’. Application of a skill $a$ at an object/region with location ${\mathbf{x}}$ at time $t\geq 0$ is denoted by $s(a,{\mathbf{x}},t)$ or, for brevity, when it is clear from the context, by $s(t)$. The time step $t$ is increased by one, once an action is completed. We assume that the robot has access to low level controllers to apply the skills in ${\mathcal{A}}$.
Partially Known Semantic Environment: The robot operates within a semantic environment $\Omega\subseteq\mathbb{R}^{d}$, $d\in\{2,3\}$ with fixed, static, and potentially unknown obstacle-free space denoted by $\Omega_{\text{free}}\subseteq\Omega$. The space $\Omega_{\text{free}}$ is populated with $M>0$ static semantic objects. Each object $e$ is characterized by its location ${\mathbf{x}}_{e}$ and semantic label $o_{e}\in{\mathcal{O}}$, where ${\mathcal{O}}$ is a set collecting all possible semantic labels that objects may have (e.g., ‘bottle’ or ‘chair’). The robot is assumed to have knowledge of both the location and label of each object. Objects may also be located inside containers (e.g., drawer or fridge), with their status (open/closed) initially unknown. Also, the occupied space $\Omega\setminus\Omega_{\text{free}}$ may prevent access to certain semantic objects rendering execution of certain decisions $s(t)$ infeasible. We assume that the robot is equipped with sensors allowing it to detect obstacles and reason about the status of containers containing objects of interest. Hereafter, we denote by ${\mathcal{S}}$ a finite set collecting all decisions $s(t)$ the robot can take. This set is constructed offline using the action space ${\mathcal{A}}$ and the available objects where each action can be applied.
Mission Specification: The robot is tasked with a high-level mission with temporal and logical requirements. To formally define such tasks, we employ Linear Temporal Logic (LTL). LTL is a formal language that comprises a set of atomic propositions (AP) (i.e., Boolean variables), denoted by $\mathcal{AP}$, Boolean operators, (i.e., conjunction $\wedge$, and negation $\neg$), and temporal operators, such as always $\square$, eventually $\lozenge$, and until $\mathcal{U}$; see also Example II.4. A formal presentation of the syntax and semantics of LTL can be found in [82]. For simplicity, hereafter, we restrict our attention to co-safe LTL formulas that is a fragment of LTL that exclude the ‘always’ operator. Co-safe LTL formulas can be satisfied within a finite horizon $H$. For brevity, throughout the paper, we may use the term LTL instead of co-safe LTL. We define APs so that they are true when a sub-task expressed in natural language (NL) is satisfied, and false otherwise. Valid definitions of APs include sub-tasks that can be accomplished by a finite robot trajectory $\tau$, defined as a finite sequence of $T$ decisions selected from ${\mathcal{S}}$, i.e.,
$\tau=s(t),s(t+1),\dots,s(t+k),\dots,s(t+T-1),$ (2)
where $k\in\{0,\dots,T-1\}$, for some $T\geq 1$; see also Example II.4 and Section V-D. We call formulas constructed in this way as co-safe LTL-NL formulas. Co-safe LTL-NL formulas are satisfied by finite-horizon robot trajectories $\tau_{\phi}$ defined as
$\tau_{\phi}=\tau_{1},\dots,\tau_{n},\dots,\tau_{N},$ (3)
where $\tau_{n}$ is a finite robot trajectory of horizon $T_{n}$, as defined in (2). Thus, the total horizon $H$ of the plan $\tau_{\phi}$ is $H=\sum_{n=1}^{N}T_{n}$. We highlight that in $\tau_{\phi}$, the index $n$ is different from the time instants $t\in\{1,\dots,H\}$. In fact, $n\in\{1,\dots,N\}$ is an index, initialized as $n=1$ and increased by $1$ every $T_{n}$ time instants, pointing to the next finite trajectory in $\tau_{\phi}$.
Our goal is to design plans $\tau_{\phi}$ satisfying co-safe LTL-NL tasks under the following assumptions:
Assumption II.1 (Error-free Skills)
The robot has access to low-level controllers allowing it to apply skills error-free.
Assumption II.2 (Independent APs)
The APs are independent of each other, meaning that the satisfaction of one AP does not prohibit the satisfaction of another AP.
Remark II.3 (Assumptions)
We make Assumption II.1 as we consider robots with known and deterministic system dynamics as in (1). This assumption will be used to provide mission completion guarantees in Section IV; see Theorem IV.2. Formally relaxing this assumption and extending the proposed planner to account for uncertainty in the system dynamics and the performance of the robot skills is part of our future work; see also Remark IV.3. Assumption II.2 is made only for simplicity as it will be discussed in Remark III.1.
Problem Statement: This paper addresses the following problem (see Ex. II.4):
Problem 1
Given a robot with capabilities ${\mathcal{A}}$, a partially unknown semantic environment $\Omega$, and a co-safe LTL-NL task $\phi$, design a robot plan $\tau_{\phi}$ satisfying $\phi$.
Example II.4
[…]
III Hierarchical Temporal Logic Planning with Natural Language Instructions
In this section, we propose HERACLEs, a hierarchical planner to address Problem 1. In Section III-A, we present an overview HERACLEs; see Alg. 1. A detailed description of its components is provided in Sections III-B-III-E.
Algorithm 1 HERACLEs: A Hierarchical Neuro-Symbolic Planner
1:Input: LTL-NL Task $\phi$; Coverage level $\alpha$;
2:Symbolic planner generates sub-task: $\pi_{\text{next}}$ and $\Sigma^{\text{unsafe}}$
3:Convert sub-task into an initial prompt $\ell(1)$ with empty history of action
4:while ($\phi$ not accomplished) $\wedge~{}(\pi_{\text{next}}\neq\varnothing)$ do
5:     $\tau=\varnothing$
6:     for $k=0~{}\text{to}~{}T-1$ do
7:         Compute the prediction set $\mathcal{C}(\ell(t+k))$ (Alg. 2)
8:         if  $|\mathcal{C}(\ell(t+k))|>1$  then
9:              Request a new sub-task ($\pi_{\text{next}}$, $\Sigma^{\text{unsafe}}$)
10:              if $\pi_{\text{next}}\neq\varnothing$ then
11:                  $t=t+k$, and go to line 5
12:              else
13:                  Obtain $s(t+k)$ from human operator
14:         else
15:              Pick (unique) decision $s(t+k)\in\mathcal{C}(\ell(t+k))$
16:         if $\Omega$ prevents execution of $s(t+k)$ then
17:              Request a new sub-task ($\pi_{\text{next}}$, $\Sigma^{\text{unsafe}}$)
18:              if $\pi_{\text{next}}\neq\varnothing$ then
19:                  $t=t+k$, and go to line 5
20:              else
21:                  Report mission failure (termination)
22:         else
23:              Execute decision $s(t+k)$
24:              Observe environment update $p(t+k)$
25:         Update $\ell(t+k+1)=\ell(t+k)+s(t+k)+p(t+k)$
26:     Construct $\tau=[s(t),\dots,s(t+T-1)]$
27:     Append $\tau$ to the plan $\tau_{\phi}$
28:     Current time step: $t=t+T$
29:     Request a new sub-task ($\pi_{\text{next}}$, $\Sigma^{\text{unsafe}}$) and check for mission accomplishment
III-A Overview of HERACLEs
HERACLEs takes as input an LTL-NL mission defined over a partially known semantic environment. This LTL-NL formula is processed online by an existing symbolic temporal logic planner that, given the current mission status, determines the next language-based sub-task the robot should accomplish to make mission progress. This symbolic planner is presented in Section III-B and is adopted from [19, 20]; we note that any other temporal logic planner can be used. This language-based sub-task serves as an input to a pre-trained LLM that is responsible for generating a feasible plan of the form (2); see Section III-C. The LLM planner communicates back to the LTL planner with one of the following messages: (i) The LLM is ‘uncertain’ about what the correct plan is [line 7-8, Alg. 1]. In this case, the LTL planner generates an alternative sub-task to proceed [lines 9-11, Alg. 1]. If such a sub-task does not exist, help from users is requested who either provide the correct decision or decide to halt operation [line 13, Alg. 1]. (ii) A plan has been designed but the (initially unknown) geometric structure of the environment ends up preventing the robot from fully executing it [line 16, Alg. 1]. In this case, the LTL planner generates an alternative sub-task to proceed and the above process repeats [lines 17-19, Alg. 1]. If such a sub-task does not exist, the mission terminates unsuccessfully (e.g., the LTL-NL mission may be infeasible) [line 21, Alg. 1]. (iii) A plan has been designed and executed, accomplishing the assigned sub-task [lines 22-24, Alg. 1]. In this case, the LTL planner generates a new sub-task, if the overall mission has not been accomplished yet, and the above process repeats [line 29, Alg. 1]. A detailed description of when the LLM requests assistance (cases (i)-(ii)) from the symbolic planner is provided in Section III-D. Conformal prediction, outlined in Section III-E, is employed for uncertainty quantification in case (i).
III-B Symbolic Temporal Logic Task Planner: What Language-based Sub-Task to Accomplish Next?
In this section, we provide an overview of an existing temporal logic task planner that we employ to determine what language sub-task the robot should accomplish next to make progress towards accomplishing the LTL-NL task $\phi$. A detailed presentation of this planner can be found in [19, 20].
Consider an LTL-NL specification $\phi$. Initially, we translate $\phi$ into a Deterministic Finite state Automaton (DFA). This step occurs offline. The DFA can be conceptualized as a graph structure, with state-space (set of nodes) and transitions (set of edges) between states/nodes. The DFA has an initial node representing the mission’s commencement, a set of final nodes denoting mission completion, and intermediate nodes representing various stages of the mission. Transitions between nodes occur when specific atomic predicates are satisfied. Mission accomplishment is achieved upon reaching a final state starting from the initial one.
At any time $t\geq 0$, the task planner receives three inputs: (i) an LTL-NL task $\phi$; (ii) the current mission status; and (iii) a set $\Sigma^{\text{unc}}(t)$ of NL-based APs/sub-tasks. The input (ii) is represented by the DFA state that has been reached, starting from the initial one, given the sequence of actions that the robot has applied up to time $t$. The set in (iii) is initially empty, i.e., $\Sigma^{\text{unc}}(0)=\emptyset$, and is dynamically updated as the robot navigates the world. Specifically, it is updated so that it collects all NL-based APs that the robot is uncertain about how to accomplish or that are physically impossible to achieve (e.g., they require reaching inaccessible objects). The details of its update are discussed in Section III-C.
Given these inputs, the task planner generates: (i) an NL-based AP, denoted by $\pi_{\text{next}}$; and (ii) a set of NL-based APs collected in a set $\Sigma^{\text{unsafe}}$ [line 2, Alg. 1]. These sets impose two conditions that must be met to make mission progress (i.e., to move closer to a final DFA node). Specifically, the robot should design a plan $\tau$, defined as in (2), to (a) accomplish the language-based sub-task captured in $\pi_{\text{next}}$ while, (b) in the meantime, not satisfying any of the APs collected in $\Sigma^{\text{unsafe}}$. Note that the set $\Sigma^{\text{unsafe}}$ depends on the current mission status and, therefore, it does not remain fixed throughout the mission. These two requirements act as inputs to an LLM-based planner; refer to Section III-C. The LLM-based planner is responsible for designing a plan $\tau$ that satisfies (a)-(b).
Remark III.1 (Relaxing Assumption II.2)
The temporal logic planner used in Section III-B is myopic as it computes the next NL-based AP $\pi_{\text{next}}$ for the robot, instead of a sequence of sub-tasks that need to be accomplished to satisfy the overall mission. Assumption II.2 is a necessary condition required to ensure completeness of that symbolic planner. This assumption can be relaxed by employing non-myopic symbolic planners, which generate a sequence of predicates/sub-tasks (rather than only the next sub-task) that the robot should accomplish to complete the mission [14, 28]. The benefit of using the employed myopic symbolic planner, as opposed to a non-myopic one, lies in its computational efficiency, as demonstrated in [19, 20].
III-C LLM-based Planner: How to Accomplish the Assigned Language-based Sub-Task?
To synthesize a finite horizon plan $\tau$ satisfying conditions (a)-(b), we employ pre-trained LLMs. Hereafter, we simply refer to the requirements (a)-(b) as the (language-based) sub-task that the robot should accomplish next. We convert this planning problem into a sequence of $T>0$ multiple-choice question-answering (MCQA) problems for the LLM where $T$ is a hyperparameter. [^1] Essentially, $\tau$ and $T$ refer to $\tau_{n}$ and $T_{n}$ defined in Section II. For ease of presentation, we drop the dependence on the index $n$. The ‘question’ refers to the sub-task along with any actions that the robot has taken up to time $t+k$, $k\in\{0,\dots,T-1\}$ to accomplish it. We denote by $\ell(t+k)$ the (textual) description of the ‘question’. The ‘choices’ refer to the decisions $s(t+k)$ that the robot can make. Given $\ell(t+k)$, the LLM will select $s(t+k)$. This process occurs sequentially for $k\in\{0,\dots,T-1\}$ giving rise to a plan $\tau$. In what follows, we discuss how $\ell(t+k)$ is structured and then we show how it can be used to compute $s(t+k)$.
Figure 2: Example of a prompt constructed in our simulations in Section V. This prompt refers to $t=1$ when the history of actions is empty.
Prompt Construction: In what follows, we discuss how we structure the context text (‘prompt’) that will be given as an input to the LLM; see also Fig. 2. The prompt used in this work consists of the following parts. (A) System description that defines the action space determining all possible actions $a\in{\mathcal{A}}$ that the robot can apply as well as the objective of the LLM. (B) Environment description that describes the locations ${\mathbf{x}}_{e}$ of each semantic object $e$ of interest; (C) Task description that includes the language-based task $\pi_{\text{next}}$ (condition (a)) as well as language-based constraints (if any), modeled by $\Sigma^{\text{unsafe}}$ that the robot should respect until $\pi_{\text{next}}$ is satisfied (condition (b)); (D) History of actions & current environment status that includes the sequence of actions, generated by the LLM, that the robot has executed so far towards accomplishing the assigned task. It also includes the current locations of semantic objects that the robot may have manipulated/moved so far; (E) Response structure describing the desired structure of the LLM output for an example task.
Plan Design & Execution: At iteration $k=0$, part (D) does not include any textual information as a new sub-task has just been announced [line 3, alg. 1]. Given $\ell(t+k)$, the LLM is asked to make a decision $s(t+k)$ among all available ones included in part (A). The LLM selects $s(t+k)$ as follows. Given any $s\in{\mathcal{S}}$, LLMs can provide a score $g(s|\ell(t+k))$; the higher the score, the more likely the decision $s$ is a valid next step to address the language-instruction provided in $\ell(t+k)$. To get $g(s|\ell(t+k))$, we query the LLM over all potential decisions $s\in{\mathcal{S}}$ [50].
Using these scores, a possible approach to select $s(t+k)$ is by simply choosing the decision with the highest score, i.e., $s(t+k)=\arg\max_{s\in\mathcal{S}}g(s|\ell(t+k))$. However, these scores do not represent calibrated confidence. Thus, we calibrate these confidence scores and let the LLM make decisions only when it is certain enough [41]. We formalize this by constructing a set of actions (called, hereafter, prediction set), denoted by ${\mathcal{C}}(\ell(t+k))$, that contain the ground truth action with user-specified confidence [line 7, alg. 1]. Hereafter, we assume that such prediction sets are provided; we defer their construction, using conformal prediction, to Section III-E (see (12)). Given ${\mathcal{C}}(\ell(t+k))$, we select the decision $s(t+k)$ as follows. If $|{\mathcal{C}}(\ell(t+k))|=1$, then we select the decision included in ${\mathcal{C}}(\ell(t+k))$ as it contributes to mission progress with high confidence [line 15, Alg. 1]. [^2] By construction of the prediction sets, this action coincides with $s(t+k)=\arg\max_{s\in\mathcal{S}}g(s|\ell(t+k))$; see Section III-E. As soon as it is selected, the robot physically executes it using its library of low-level controllers [line 23, alg. 1]. If $|{\mathcal{C}}(\ell(t+k))|>1$ [line 8-13, alg.1] or if $|{\mathcal{C}}(\ell(t+k))|=1$ but the robot cannot physically execute $s(t+k)$ the robot seeks assistance to select $s(t+k)$ [line 16-21, Alg.1]. The process of asking for assistance is described in more detail in Section III-D.
Once a decision $s(t+k)$ is made and physically executed, the current time step is updated to $t+k+1$. Then, the prompt $\ell(t+k+1)$ is constructed that will be used to select $s(t+k+1)$. Parts (A)-(C) and (E) in the prompts $\ell(t+k)$ and $\ell(t+k+1)$ are the same. Part (D) in $\ell(t+k+1)$ is augmented by recording the decision $s(t+k)$ as well as incorporating perceptual feedback about the status of containers (if any) that may contain any of the semantic objects. We automatically convert this perceptual feedback into text denoted by $p(t+1)$ (see Fig. 2) [line 24, Alg. 1]. With slight abuse of notation, we denote this prompt update by
$h(t+1)=h(t)+s(t)+p(t+1),$
where the summation means concatenation of text [line 25, Alg.1]. This process is repeated for all $k\in\{0,\dots,T-1\}$ to sequentially select $s(t+k)$. This process generates a plan $\tau$ of length $T$ as in (2) [line 26, alg. 1]. At time $t+T$, part (D) in $\ell(t+T)$ is updated so that it does not include any information since a new sub-task will be announced by the LTL task planner [line 29, alg. 1]. Concatenation of all plans $\tau$ for the sub-tasks generated by the LTL planner gives rise to the plan $\tau_{\phi}$ [line 27, alg. 1] (see Section II).
III-D When to Seek Assistance?
Assume that there exists at least one $k\in\{0,\dots,T-1\}$ so that $|\mathcal{C}(\ell(t+k)))|>1$. In this case, the robot asks for help in order to proceed [line 8-13, Alg. 1]. This assistance request and response occurs as follows. Initially, the robot requests from the LTL task planner an alternative sub-task to make mission progress. To make this help-request, the robot updates the set $\Sigma^{\text{unc}}$, defined in Section III-B, as $\Sigma^{\text{unc}}=\Sigma^{\text{unc}}\cup\{\pi_{\text{next}}\}$. The resulting set $\Sigma^{\text{unc}}$ along with the current mission status are sent to the LTL task planner which then generates a new sub-task [line 9, alg. 1]. If such a sub-task exists, the process discussed in Section III-C repeats [line 11, alg. 1]. If there are no alternative sub-tasks to proceed, then the LTL planner fails to provide assistance. In this case, the robot asks for help from a human operator [line 13, alg. 1]. Specifically, the robot returns to the user the current prompt $\ell(t+k)$ along with the prediction set $\mathcal{C}(\ell(t+k)))$ for which it holds $|\mathcal{C}(\ell(t+k)))|>1$. The user selects the correct decision $s(t+k)$ if it exists in the prediction set or it halts operation. Help from the LTL planner is also requested if the robot cannot physically execute the selected decision $s(t+k)$. This can occur if the robot detects that the initially unknown geometric structure of the environment prevents it from reaching desired destinations/objects, as in [19, 20] [line 16-21, Alg. 1]; see Sec. V-C. In this case, if the LTL planner cannot generate alternative sub-tasks the mission terminates unsuccessfully [line 21, alg. 1].
III-E Constructing Prediction Sets using Conformal Prediction
Algorithm 2 Computation of Causal Prediction Sets $\mathcal{C}({\ell}_{\text{test}}(t))$
1:Inputs: Calibration set ${\mathcal{M}}$; New mission scenario; Current prompt ${\ell}_{\text{test}}(t)$; Threshold $\alpha$
2:Output: Causal Prediction Set $\mathcal{C}({\ell}_{\text{test}}(t))$
3:if $t=1$ then
4:     Compute $\bar{r}_{i}$ for all calibration sequences $i$ (see (9))
5:     Compute the $\frac{(M+1)(1-\alpha)}{M}$ quantile $\bar{q}$ of $\{\bar{r}_{i}\}_{i=1}^{M}$
6: Compute the LLM confidence $g(s|{\ell}_{\text{test}}(t))$, for each $s\in{\mathcal{S}}$
7: Return the prediction set $\mathcal{C}({\ell}_{\text{test}}(t))=\{s~{}|~{}g(s|{\ell}_{\text{test}}(t))>1-\bar{q}\}$ (see (13)).
In this section, we discuss how the prediction sets ${\mathcal{C}}(\ell(t))$, introduced in Section III-C, are constructed. To construct them, we employ conformal prediction (CP), a statistical tool for uncertainty quantification in black box models [65]. Our analysis builds upon [41, 60]. Construction of these prediction sets requires a distribution ${\mathcal{D}}$ from which LTL-NL missions are sampled. This is important as the application of CP requires the construction of a calibration dataset collecting missions, along with the corresponding ground truth plans, that are independently generated by ${\mathcal{D}}$. Then, the prediction sets for a new/validation mission are ‘correct’ as long as the calibration and the validation missions are i.i.d.; a discussion on relaxing this i.i.d. assumption is provided in Remark IV.4.
[…]
Single-step Plans: We begin by considering LTL-NL formulas $\phi$ that can be satisfied by plans $\tau_{\phi}$ of horizon $H=1$ (see Section II); later we generalize the results for $H\geq 1$. This also means that synthesizing $\tau_{\phi}$ requires the LLM to make a single decision $s$. First, we sample $M$ independent scenarios from ${\mathcal{D}}$. We refer to these scenarios as calibration scenarios. For each calibration scenario $i\in\{1,\dots,M\}$, we construct its equivalent prompt $\ell_{\text{calib}}^{i}$. For each prompt, we (manually) compute the ground truth plan $\tau_{\text{calib}}^{i}=s_{\text{calib}}^{i}(1)$ accomplishing this task. For simplicity, we assume that there exists a unique correct decision $s_{\text{calib}}^{i}$ for each calibration scenario. [^3] This assumption can be relaxed as in [41] yielding prediction sets that contain the correct choice with the highest LLM confidence value (among all other correct choices). Hereafter, we drop the dependence on the robot decisions and prompts on the time step, since we consider single-step plans. This way we construct a calibration dataset =${\mathcal{M}}=\{(\ell_{\text{calib}}^{i},\tau_{\text{calib}}^{i})\}_{i=1}^{M}$.
Consider a new scenario drawn from ${\mathcal{D}}$, called validation/test scenario. We convert this scenario into its equivalent prompt $\ell_{\text{test}}$. Since the calibration and the validation scenario are i.i.d., CP can generate a prediction set ${\mathcal{C}}(\ell_{\text{test}})$ of decisions $s$ containing the correct one $s_{\text{test}}$ with probability greater than $1-\alpha$, i.e.,
$P(s_{\text{test}}\in\mathcal{C}(\ell_{\text{test}}))\geq 1-\alpha,$ (4)
where $\alpha\in[0,1]$ is user-specified. To generate $\mathcal{C}(\ell_{\text{test}})$, CP first uses the LLM’s confidence score $g$ (see Section III-C) to compute the set of non-conformity (NC) scores $\{r_{i}=1-g(s_{\text{calib}}^{i}~{}|~{}\ell_{\text{calib}}^{i})\}_{i=1}^{M}$ over the calibration set. The higher the NC score $r_{i}$ is, the worse the performance of the LLM is at the $i$ calibration point. Then CP performs calibration by computing the $\frac{(M+1)(1-\alpha)}{M}$ empirical quantile of $r_{1},\dots,r_{M}$ denoted by $q$. Then, it generates the prediction set
$\mathcal{C}(\ell_{\text{test}})=\{s\in{\mathcal{S}}~{}|~{}g(s|\ell_{\text{test}})>1-q\},$ (5)
that includes all decisions that the predictor is at least $1-q$ confident in. The generated prediction set ensures the $1-\alpha$ marginal coverage guarantee in (4) holds. [^4] To obtain a meaningful prediction set, it should hold $M\geq\big{\lceil}(M+1)(1-\alpha)\big{\rceil}$; otherwise the prediction set will be empty. This coverage guarantee is marginal in the sense that the probability is defined over the randomness of the calibration dataset and the validation scenario. By construction of the prediction sets the decision $s=\arg\max_{s\in\mathcal{S}}g(s|\ell_{\text{test}})$ belongs to $\mathcal{C}(\ell_{\text{test}})$.
Multi-step Plans: Next, we generalize the above result to the case where satisfaction of $\phi$ requires plans $\tau_{\phi}$ with $H\geq 1$ decisions selected from ${\mathcal{S}}$; see Alg. 2. Here we cannot apply directly (4)-(5) to compute individual sets ${\mathcal{C}}(\ell_{\text{test}}(t))$ for the robot, as this violates the i.i.d. assumption required to apply CP. The challenge in this case is that the prompts $\{(\ell_{\text{test}}(t)\}_{t=1}^{H}$ are not independent of each other which violates the i.i.d. assumption required to apply CP. In fact these prompts depend on past robot decisions as well as on the LTL-NL tasks $\phi_{\text{test}}$. To address this challenge, inspired by [41], we (i) lift the data to sequences, and (ii) perform calibration at the sequence level using a carefully designed NC score function.
First, we construct a calibration dataset as follows. We generate $M\geq 1$ scenarios $\xi_{i}$ from ${\mathcal{D}}$. The LTL-NL formula $\phi_{i}$ of each scenario is broken into a sequence of $H_{i}\geq 1$ prompts, defined as: [^5] The distribution ${\mathcal{D}}$ over scenarios induces a distribution over data sequences (6) [41]. These data sequences are equivalent representations of the sampled scenarios augmented with the ground truth decisions.
¯$\bar{\ell}_{\text{calib}}^{i}=[{\ell}_{\text{calib}}^{i}(1),\dots,{\ell}_{\text{calib}}^{i}(H_{i})],$ (6)
where each prompt in the sequence ¯$\bar{\ell}_{\text{calib}}^{i}$ contains a history of ground truth decisions made so far. We define the corresponding sequence of ground truth decisions as:
$\tau_{\phi,\text{calib}}^{i}=s_{\text{calib}}^{i}(1),\dots,s_{\text{calib}}^{i}(H_{i}),$ (7)
This gives rise to the calibration set ={(¯${\mathcal{M}}=\{(\bar{\ell}_{\text{calib}}^{i},\tau_{\phi,\text{calib}}^{i})\}_{i=1}^{M}$. As before, we assume that each context ¯$\bar{\ell}_{\text{calib}}^{i}$ has a unique correct plan $\tau_{\phi,\text{calib}}^{i}$. Next, we use the lowest score $g$ over the time-steps $1,\dots,H_{i}$ as the score for each sequence $i$ in calibration set, i.e.,
$\bar{g}(\tau_{\phi,\text{calib}}^{i}~{}|~{}\bar{\ell}_{\text{calib}}^{i})=\min_{t\in\{1,\dots,H_{i}\}}g(\tau_{\phi,\text{calib}}^{i}(t)~{}|~{}{\ell}_{\text{calib}}^{i}(t)).$ (8)
Thus, the NC score of each sequence $i$ is [line 4, Alg. 2]
$\bar{r}_{i}=1-\bar{g}(\tau_{\phi,\text{calib}}^{i}|\bar{\ell}_{\text{calib}}^{i}).$ (9)
Consider a new scenario $\xi_{\text{test}}$ associated with a task $\phi_{\text{test}}$ with horizon $H_{\text{test}}$. This scenario corresponds to a sequence of prompts
¯$\bar{\ell}_{\text{test}}={\ell}_{\text{test}}(1),\dots,{\ell}_{\text{test}}(k),\dots,{\ell}_{\text{test}}(H_{\text{test}}).$ (10)
CP can generate a prediction set $\bar{{\mathcal{C}}}(\bar{\ell}_{\text{test}})$ of plans $\tau_{\phi}$ containing the correct one $\tau_{\phi,\text{test}}$ with high probability i.e.,
$P(\tau_{\phi,\text{test}}\in\bar{\mathcal{C}}(\bar{\ell}_{\text{test}}))\geq 1-\alpha,$ (11)
where the prediction set $\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})$ is defined as
$\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})=\{\tau_{\phi}~{}|~{}\bar{g}(\tau_{\phi}|\bar{\ell}_{\text{test}})>1-\bar{q}\},$ (12)
where $\bar{q}$ is the $\frac{(M+1)(1-\alpha)}{M}$ empirical quantile of $\bar{r}_{1},\dots,\bar{r}_{M}$ [line 5, Alg. 2]. The generated prediction set ensures that the coverage guarantee in (11) holds. By construction of the prediction sets, the plan $\tau_{\phi}$ generated by the LLM belongs to $\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})$.
Causal Construction of the Prediction Set: Notice that $\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})$ is constructed after the entire sequence ¯$\bar{\ell}_{\text{test}}={\ell}_{\text{test}}(1),\dots,{\ell}_{\text{test}}(H_{\text{test}})$ is obtained. However, at every (test) time $t\in\{1,\dots,H_{\text{test}}\}$, the robot observes only the prompt ${\ell}_{\text{test}}(t)$ and not the whole sequence. In what follows, we construct the prediction set in a causal manner using only the current and past information. Specifically, at every time step $t$, we construct the local prediction set as [line 6-7, Alg. 2]
$\mathcal{C}({\ell}_{\text{test}}(t))=\{s~{}|~{}g(s|{\ell}_{\text{test}}(t))>1-\bar{q}\}.$ (13)
Then, the causal prediction set for ¯$\bar{\ell}_{\text{test}}$ is defined as
$\mathcal{C}(\bar{\ell}_{\text{test}})=\mathcal{C}({\ell}_{\text{test}}(1))\times\mathcal{C}({\ell}_{\text{test}}(2))\times\dots\times\mathcal{C}({\ell}_{\text{test}}(H_{\text{test}})).$ (14)
In Section IV, we show that $\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})=\mathcal{C}(\bar{\ell}_{\text{test}})$.
Remark III.2 (Dataset-Conditional Guarantee)
The probabilistic guarantee in (11) is marginal in the sense that the probability is over both the sampling of the calibration set ${\mathcal{M}}$ and the validation point ¯$\bar{\ell}_{\text{test}}$. Thus, a new calibration set will be needed for every single test data point ¯$\bar{\ell}_{\text{test}}$ to ensure the desired coverage level. A dataset-conditional guarantee which holds for a fixed calibration set can also be applied [83, 41].
IV Probabilistic Task Satisfaction Guarantees
In this section, we show that given any (validation) scenario $\xi_{\text{test}}=\{{\mathcal{A}}_{\text{test}},\phi_{\text{test}},H_{\text{test}},\Omega_{\text{test}}\}$ drawn from a distribution ${\mathcal{D}}$, the probability that HERACLEs will design a plan satisfying the LTL-NL formula is at least $1-\alpha$, where $\alpha$ is the coverage level used to construct the prediction sets. To show this, we need first to show the following result. The proofs follow a similar logic to the ones in [41].
Proposition IV.1
The prediction set $\bar{{\mathcal{C}}}(\bar{\ell}_{\text{test}})$ defined in (12) is the same as the on-the-fly constructed prediction set ${\mathcal{C}}(\bar{\ell}_{\text{test}})$ defined in (14), i.e., $\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})=\mathcal{C}(\bar{\ell}_{\text{test}})$.
Proof:
It suffices to show that if $\tau_{\phi}$ belongs to $\bar{{\mathcal{C}}}(\bar{\ell}_{\text{test}})$ then it also belongs to ${{\mathcal{C}}}(\bar{\ell}_{\text{test}})$ and vice-versa. First, we show that if $\tau_{\phi}\in\bar{{\mathcal{C}}}(\bar{\ell}_{\text{test}})$ then $\tau_{\phi}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$. Since $\tau_{\phi}\in\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})$, then we have that $\min_{t\in\{1,\dots,H_{\text{test}}\}}g(\tau_{\phi}(t)|{\ell}_{\text{test}}(t))>1-\bar{q}$ due to (8). This means that the score $g(\tau_{\phi}(t)|{\ell}_{\text{test}}(t))>1-\bar{q}$, for all $t\in\{1,\dots,H_{\text{test}}\}$. Thus, $\tau_{\phi}(t)\in{\mathcal{C}}({\ell}_{\text{test}}(t))$, for all $t\in\{1,\dots,H_{\text{test}}\}$. By definition of ${\mathcal{C}}(\bar{\ell}_{\text{test}})$ in (14), this implies that $\tau_{\phi}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$. These steps hold in the other direction too showing that if $\tau_{\phi}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$ then $\tau_{\phi}\in\bar{{\mathcal{C}}}(\bar{\ell}_{\text{test}})$. ∎
Theorem IV.2
Consider test mission scenarios $\xi_{\text{test}}$ drawn from ${\mathcal{D}}$. Assume that prediction sets are constructed causally with coverage level $1-\alpha$ and that the robot seeks help from a user whenever the local prediction set ${\mathcal{C}}({\ell}_{\text{test}}(t))$, defined in (13), is not singleton; see Section III-D. If the LTL task planning algorithm is complete and Assumption II.1 holds, then the completion rate over new test scenarios (and the randomness of the calibration sets) drawn from ${\mathcal{D}}$ is at least $1-\alpha$ [^6] Theorem IV.2 implies that HERACLEs can achieve $1-\alpha$ mission success rates even if alternative symbolic planners are used as long as they are complete.
Proof:
If the symbolic planning algorithm is complete, then this means if there exists a solution it will find it. By solution, here we refer to a sequence of sub-tasks in the unknown environment that, if completed, the LTL mission will be satisfied. Since the scenarios $\xi_{\text{test}}$ are drawn from ${\mathcal{D}}$, this means that they are feasible by assumption (see Section III-E). This equivalently means that if the symbolic task planning algorithm is complete, any failures of Alg. 1 in finding a correct plan are not attributed to the symbolic planner. Under this setting, the following three cases may occur as the robot designs its plan. Case I: We have that $|\mathcal{C}({\ell}_{\text{test}}(t))|=1$, $\forall t\in\{1,\dots,H_{\text{test}}\}$ (possibly after asking for help from the symbolic planner) and $\tau_{\phi,\text{test}}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$ where $\tau_{\phi,\text{test}}$ is the ground truth plan and prediction set ${\mathcal{C}}(\bar{\ell}_{\text{test}})$ is defined as in (14). In this case, the robot will select the correct plan. Case II: We have that $|\mathcal{C}({\ell}_{\text{test}}(t))|>1$ (even after asking for help from the symbolic planner), for at least one time step $t\in\{1,\dots,H_{\text{test}}\}$ and $\tau_{\phi,\text{test}}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$. In this case, the robots will select the correct plan assuming users who faithfully provide help. Case III: We have that $\tau_{\phi,\text{test}}\notin{\mathcal{C}}(\bar{\ell}_{\text{test}})$. The latter means that there exists at least one time step $t$ such that $\tau_{\phi,\text{test}}(t)\notin{\mathcal{C}}({\ell}_{\text{test}}(t))$. In this case, the robot will compute an incorrect plan. Notice that we do not know which of these three cases will occur at each step $t$, since the ground truth plan for the test mission scenario is unknown. However, the probability that either Case I or II will occur is equivalent to the probability $P(\tau_{\phi,\text{test}}\in{\mathcal{C}}(\bar{\ell}_{\text{test}}))$. Due to Proposition IV.1 and (11), we have that $P(\tau_{\phi,\text{test}}\in{\mathcal{C}}(\bar{\ell}_{\text{test}}))\geq 1-\alpha$. Thus, either of Case I and II will occur with probability that is at least equal to $1-\alpha$. Since Cases I-III are mutually and collectively exhaustive, we conclude that the probability that Case III will occur is less than $\alpha$. This means that the probability of HERACLEs generating a correct plan is at least $1-\alpha$. Due to Assumption II.1, we also note that the mission success rate is at least $1-\alpha$ completing the proof. ∎
Remark IV.3 (Relaxing Assumption II.1)
Assumption II.1 requires perfect execution of robot skills and is required to provide the mission completion guarantees discussed in Theorem IV.2. Relaxing this assumption necessitates integrating skill imperfections in the conformal prediction component. Construction of such prediction sets requires developing score functions that (heuristically) quantify the likelihood of success of a given skill from the current state. A potential approach to address this is by leveraging the affordance functions learned in [50]. Selecting robot states that maximize these affordance functions can be accomplished by leveraging action-oriented semantic maps [84]. Subsequently, we can compute the product of this skill score function with the score function that heuristically captures LLM uncertainty (modeled by the LLM confidence value). This way, a higher product score indicates a greater likelihood that the corresponding action is correct and will be executed successfully. Using this product score function, CP can be applied as in Section III-E to construct prediction sets. In cases of non-singleton sets, among all correct actions (if any) that are included in the prediction set, the user will select the one with the highest likelihood of successful execution. Formally relaxing this assumption and designing LLM-based planners capable of reasoning about uncertainty in robot skills is part of our future work.
Remark IV.4 (Relaxing i.i.d. Assumption)
Theorem IV.2 holds under the assumption that validation and test scenarios are sampled independently from a distribution ${\mathcal{D}}$. We emphasize that this a common requirement in related works that employ CP for uncertainty quantification; see e.g., the works in [65, 60, 41, 66, 67, 68, 69, 70]. This assumption can be relaxed by employing robust CP to obtain valid prediction sets for all distributions ${\mathcal{D}}^{\prime}$ that are ‘close’ to $D$ (based on the flimit-$f-$divergence) [85]. Integrating HERACLEs with robust CP is out of the scope of this work.
V Numerical Experiments
In this section, we provide extensive comparative experiments against existing LLM planners that require the task description in NL. In Section V-A, we compare HERACLEs against baselines that do not allow robots to ask for help [50, 81, 40]. These experiments show that the performance gap between these methods and HERACLEs increases significantly as task complexity increases even when the help-mode is disabled from our planner. In Section V-B, we compare our planner against KnowNo [41], a baseline that also employs CP allowing robots to ask for help in cases of high uncertainty. Our comparisons show that HERACLEs achieves significantly lower help rates. In that section, we also empirically validate the theoretical mission success rate guarantees discussed in Section IV. In Section V-C, we illustrate the ask-for-help mode of our planner on mobile manipulation tasks; see also [86]. In Section V-D, we demonstrate how various definitions of the NL-based predicates may affect performance of HERACLEs. Finally, in Section V-E, we provide examples to compare LTL and LTL-NL formulas in terms of their user-friendliness. In all case studies, we pick GPT-3.5, Llama 2-13b, and Llama 3-8b as the LLMs.
Model Method Case Study
I (Easy) II (Medium) III (Hard)
GPT 3.5 Ours $96\%$ $93.3\%$ $93\%$
SayCan [50] $96\%$ $40\%$ $14.08\%$
CMAS [81] $56\%$ $20\%$ $0\%$
ProgPrompt [40] $48\%$ $16\%$ $2.5\%$
Llama 2-13b Ours $86\%$ $81\%$ $80\%$
SayCan [50] $86\%$ $18\%$ $10\%$
CMAS [81] $53\%$ $9\%$ $0\%$
ProgPrompt [40] $46.7\%$ $9\%$ $0\%$
Llama 3-8b Ours $87.5\%$ $90.9\%$ $90\%$
SayCan [50] $87.5\%$ $36\%$ $18\%$
CMAS [81] $50\%$ $27\%$ $18\%$
ProgPrompt [40] $43.8\%$ $18.1\%$ $0\%$
TABLE I: Comparison of planning accuracy of HERACLEs (when the CP/help mode is deactivated) against non-conformalized baselines, using GPT 3.5, Llama 2-13b, and Llama 3-8b.
V-A Comparisons against Non-Conformalized Neural Planners
[…]
Figure 3: Case Study I: Execution of a robot plan for the LTL-NL mission: $\Diamond$ (’Move Coke 1 to Location B’) requiring the robot to eventually deliver Coke 1 to location B. The left snapshot illustrates the robot plan generated by HERACLEs. The right snapshot shows the environment when the task is completed.
Figure 4: Case Study II: Execution of a robot plan for the LTL-NL mission: $\phi=\Diamond\pi_{1}\wedge\Diamond\pi_{2}\wedge(\neg\pi_{1}{\mathcal{U}}\pi_{2})$, where $\pi_{1}$ refers to ‘Deliver Apple to A’ and $\pi_{2}$ refers to ‘Deliver Coke 2 to A’. Notice here $\phi$ requires that the apple should not be delivered to Location A prior to Coke 2. The left two snapshots illustrate the sub-tasks that the robot accomplished in the order that they were generated by the symbolic planner. The corresponding robot plans are also depicted. The right snapshot shows the environment when the task is completed.
Baselines: As the baselines for our experiments, we employ three state-of-the-art LLM-based planners: SayCan [50], CMAS [81], and ProgPrompt[40]. Unlike HERACLEs, these baselines require the overall mission to be fully described using NL. Thus, we manually convert LTL-NL tasks into NL ones, which are then used as inputs for the baselines. SayCan generates decisions, taking into consideration the likelihood that the corresponding action will be executed correctly. Since we consider perfect execution of robot actions, we assume that the affordance functions in [50] always return a value of $1$ for all decisions. SayCan converts the NL mission planning problem into a sequence of MCQA problems, as HERACLEs does too. Then SayCan selects the action with the highest softmax score: $s(t)=\arg\max_{s\in\mathcal{S}}g(s|\ell(t))$ for every MCQA problem. We also compare our method against the CMAS planner [81]. CMAS is originally developed for teams of $N$ robots. We applied it in our setting using $N=1$. Note that CMAS and ProgPrompt do not employ the MCQA setup (unlike HERACLEs and SayCan). Instead, at every time step, the LLM generates new tokens corresponding to robot actions. ProgPrompt - unlike our method, Saycan, and CMAS - requires its prompt to be written in the format of an executable programming language such as Python. We compare the planners based on their runtimes as well as their accuracy over scenarios drawn from ${\mathcal{D}}$ defined as the percentage of scenarios where a planner generates a correct plan; see Tables I-II.
To make our comparisons fair, we have enforced the following requirements: (i) All methods are exposed to the same set of actions ${\mathcal{S}}$. (ii) All methods (except ProgPrompt) share the same prompt structure. The only difference is that part (C) in the prompts for HERACLEs includes the description of the sub-task generated by the symbolic planner while part (C) for SayCan, CMAS, and ProgPrompt includes the overall mission expressed in NL (as these baselines do not consider task decomposition). The main body of the prompt in ProgPrompt is described in Python-like format as in the original work [40]. (iii) All actions are executed perfectly. (iv) We completely deactivate the help mode from our planner since this capability does not exist in these baselines. We emphasize that we have applied (iv) only to make comparisons fair against the baselines as otherwise our planner can outperform them by picking a low enough value for $\alpha$ and asking for help from users. This choice also allows us to assess the ‘nominal’ performance of our planner when help from users is not available; evaluation of our method without removing CP is presented in Sections V-B. The requirement (iv) means that we remove altogether the CP component from our planner. This implies that HERACLEs always picks the decision $s(t)=\arg\max_{s\in\mathcal{S}}g(s|\ell(t))$ (as SayCan does too). Also, to accommodate (iv), we consider environments with known obstacles that do not prevent the robot from accomplishing any sub-task included in $\phi$. This implies that HERACLEs will never trigger an assistance request from the symbolic planner. We note again that have applied (iv) only to ensure fairness to other baselines that do not have the capability of asking for help from users or symbolic planners; demonstrations in environments that prevent the accomplishment of certain sub-tasks are provided in Section V-C. (v) We require all methods to complete the plan within $H=T\times K$ steps, where $K$ is the number of predicates in $\phi$ and $T=7$ is the hyper-parameter used in Section III-C.
In what follows, we sample $111$ scenarios from ${\mathcal{D}}$ and group them based on their difficulty category. We first discuss the planning accuracy of all methods when they are coupled with GPT-3.5 and Llama 2-13b. These results are summarized in Table I which also includes the performance of the methods when equipped with Llama 3-8b.
Case Study I (Easy): We consider $25$ LTL-NL formulas of the form $\phi=\Diamond\pi_{1}$ where $\pi_{1}$ is defined as ‘Move object $e$ to location ${\mathbf{x}}$” for various objects $o_{e}\in{\mathcal{O}}$ and locations ${\mathbf{x}}$. We manually translate such formulas into NL as ‘Eventually move object $e$ to location ${\mathbf{x}}$’ to use as the task specification for CMAS and SayCan. Using GPT 3.5, the accuracy of HERACLEs was $96\%$ while the accuracy of SayCan, CMAS, and ProgPrompt were $96\%$, $56\%$, and $48\%$, respectively. Using Llama 2-13b, the accuracy of our method and SayCan dropped to 86%, while for CMAS and ProgPrompt, it dropped to 53% and 46.7%, respectively. This accuracy drop is expected as Llama 2-13b is a smaller model compared to GPT 3.5. Notice that the performance of our method and SayCan is the same since the mission $\phi$ cannot be further decomposed and both methods share the same prompt as well as the same action selection mechanism when the help mode is removed from our method. A robot plan generated by HERACLEs using GPT 3.5 is shown in Fig. 3.
Case Study II (Medium): We consider $15$ LTL-NL formulas defined as either $\phi_{1}=\Diamond\pi_{1}\wedge\Diamond\pi_{2}$ or $\phi_{2}=\Diamond\pi_{1}\wedge\Diamond\pi_{2}\wedge(\neg\pi_{1}{\mathcal{U}}\pi_{2})$. The task $\phi_{1}$ requires to eventually complete the sub-tasks $\pi_{1}$ and $\pi_{2}$ in any order while $\phi_{2}$ requires $\pi_{2}$ to be completed strictly before $\pi_{1}$. The APs $\pi_{1}$ and $\pi_{2}$ are defined as before. Using GPT 3.5, the accuracy of our planner, SayCan, CMAS, and ProgPrompt are $93.3\%$, $40\%$, $20\%$, and $16\%$, respectively. Using Llama 2-13b, the accuracy of our method is $81\%$, while SayCan dropped to $18\%$; CMAS and ProgPrompt both achieved only $9\%$ accuracy. Observe that the performance of the baselines drops as temporal and/or logical requirements are incorporated into the task description. A robot plan synthesized for one of the considered scenarios using HERACLEs and GPT 3.5 is shown in Fig. 4.
Case Study III (Hard): We consider $71$ LTL-NL formulas defined over four to six predicates defined as before. Two examples of such LTL formulas are: $\phi_{1}=\Diamond\pi_{1}\wedge\Diamond\pi_{2}\wedge\Diamond\pi_{3}\wedge\Diamond\pi_{4}\wedge(\neg\pi_{4}{\mathcal{U}}\pi_{1})$ and $\phi_{2}=\Diamond\pi_{1}\wedge\Diamond\pi_{2}\wedge\Diamond\pi_{3}\wedge(\neg\pi_{3}{\mathcal{U}}\pi_{2})\wedge\Diamond\pi_{5}\wedge(\neg\pi_{2}{\mathcal{U}}\pi_{5})\wedge(\neg\pi_{5}{\mathcal{U}}\pi_{1})\wedge\Diamond\pi_{4}$. For instance, $\phi_{1}$ requires the robot to accomplish $\pi_{1}$, $\pi_{2}$, $\pi_{3}$ and $\pi_{4}$ in any order as long as $\pi_{1}$ is executed before $\pi_{4}$. The accuracy of our planner, SayCan, CMAS, and ProgPrompt is $93\%$, $14.08\%$, $0\%$, and $2.5\%$, respectively, using GPT 3.5. As we repeat the tasks using Llama 2-13b, the accuracy of our method and Saycan are $80\%$ and $10\%$, respectively, while both CMAS and ProgPrompt failed to generate any feasible plan resulting in $0\%$ accuracy. Mistakes made by our planner were mostly because the LLM asked the robot to move to the wrong location to pick up a desired object or the LLM requested the robot to pick up an object inside a closed container without first opening it.
Runtimes: In the above case studies, the average runtime of the symbolic planner of HERACLEs to generate a sub-task was $2.7\times 10^{-5}$ secs; the runtime to generate the entire plan for different methods using GPT 3.5, Llama 2-13b, and Llama 3-8b is reported in Table II. Observe that HERACLEs and SayCan have comparable runtimes as they both rely on the MCQA framework for action selection, requiring $|{\mathcal{S}}|$ API calls to select $s(t)$. In contrast, CMAS and ProgPrompt require less runtime across all models because they need only one API call at each time $t$ to design $s(t)$. Additionally, note that the runtimes for all methods are significantly higher when using Llama 2-13b and Llama 3-8b models compared to GPT-3.5. This is because Llama models are stored and run on our local computers, whereas GPT-3.5 is accessed from OpenAI servers.
Summary of Comparisons: Observe that the performance gap, in terms of planning accuracy, between HERACLEs and the baselines, increases significantly as the task complexity increases. Also, notice that the performance of HERACLEs does not change significantly across the considered case studies. The reason is that HERACLEs decomposes the overall planning problem into smaller ones that can be handled efficiently by the LLM. This is in contrast to SayCan, CMAS, and ProgPrompt, where the LLM is responsible for generating plans directly for the original long-horizon task. Additional comparisons showing that the ability of LLMs to design correct plans deteriorates as temporal and logical requirements are incorporated into the mission can be found in [57]. Also, observe that SayCan performs better than CMAS and ProgPrompt. We attribute this empirical result to the MCQA framework, as it attempts to eliminate hallucinations. To the contrary, CMAS and ProgPrompt require the employed LLMs to generate new tokens to design robot plans increasing the risk of hallucinations.
Model Method Average Runtime (s)
GPT 3.5 Ours 14
SayCan [50] 15
CMAS[81] 3
ProgPrompt[40] 4
Llama 2-13b Ours 71
SayCan [50] 75
CMAS[81] 15
ProgPrompt[40] 18
Llama 3-8b Ours 76
SayCan [50] 78
CMAS[81] 17
ProgPrompt[40] 25
TABLE II: Comparison of average runtime for plan synthesis of HERACLES (when the CP/Help mode is deactivated) against non-conformalized baselines, using GPT 3.5, Llama 2-13b, and Llama 3-8b.
Model Method Help Rate
$1-\alpha=90\%$ $1-\alpha=95\%$
Llama 2-13B Ours $9\%$ $12\%$
KnowNo $81\%$ $87\%$
Llama 3-8B Ours $20\%$ $24\%$
KnowNo $67\%$ $71\%$
TABLE III: Comparison of help rates of HERACLEs (with CP/help mode activated) against KnowNo, a conformalized baseline, using Llama 2-13b and Llama 3-8b.
V-B Comparisons against Conformalized Neural Planners
The average planning accuracy of our method across the $111$ scenarios considered in Section V-A is $81.26\%$ and $89.7\%$ using Llama 2-13b and Llama 3-8b, respectively. In what follows, first we empirically validate the theoretical guarantees for the mission success rate of HERACLEs using both Llama 2-13b and Llama 3-8b model, i.e., we investigate if we can increase the mission success rate to a desired level $1-\alpha$ $(>81.26\%/89.7\%)$ by allowing the planner to ask for help, enabled by CP, when needed; see Alg. 1. For each validation scenario considered in Section V-A, we collect $20$ calibration sequences from ${\mathcal{D}}$. We compute a plan using Alg.1 while allowing the robot to ask for help from users when needed. We compute the ratio of how many of the corresponding $111$ generated plans are the ground truth plans with respect to their corresponding LTL-NL task $\phi$. We repeat the above process $100$ times. The average ratio across all experiments is the (empirical) mission success rate. We set $0.91-\alpha=$ 0.9. When HERACLEs is paired with Llama 2-13b and Llama 3-8b, its mission success rates are $93.3\%$ and $91\%$, validating Theorem IV.2, respectively. The help rates are $9\%$ and $20\%$ for Llama 2-13b and Llama 3-8b, respectively. Next, we set $0.951-\alpha=$ 0.95. Then, the mission success rates of our method are $97.2\%$ and $98.1\%$ when coupled with Llama 2-13b and Llama 3-8b, respectively. Also, the help rates are $12\%$ and $24\%$ for Llama 2-13b and Llama 3-8b, respectively. Observe that, as expected, the help rates increase as $1-\alpha$ increases.
Second, we compare HERACLEs against KnowNo [41] that requires the mission description solely in NL; see Table III. As discussed in Section I, KnowNo also allows robots to ask for help, enabled by CP, in cases of high uncertainty. We repeated the above process for KnowNo equipped with both Llama 2-13b and Llama 3-8b. Both KnowNo and HERACLEs were exposed to the same validation and calibration data. Our comparisons show that KnowNo requires significantly higher help rates than HERACLEs. Specifically, using Llama 2-13b, the help rates for KnowNo when $0.91-\alpha=$ 0.9 and $0.951-\alpha=$ 0.95 are $81\%$ and $87\%$, respectively. Observe that these help rates are significantly higher than the ones of HERACLEs. Similar observations are made when comparing the help rates of HERACLEs and KnowNo when they are paired with Llama 3-8b; see Table III. We attribute this performance gap to our symbolic planner that decomposes the LTL-NL missions into multiple ‘small’ NL tasks that the LLM can handle efficiently. In contrast in [41], the LLM is responsible for computing plans for the original mission.
Remark V.1 (LLM model & Help Rates)
The quality of the LLM model can critically affect the help rates. More effective LLM models may result in lower help rates. For instance, the help rate of our method when coupled with GPT 3.5 is $2.56\%$, $2.63\%$, and $2.70\%$ when $0.951-\alpha=$ 0.95, $0.961-\alpha=$ 0.96, and $0.971-\alpha=$ 0.97.
V-C Asking for Help from Symbolic Planner and Users
In this section, we demonstrate HERACLEs, using ROS/Gazebo [87], on mobile manipulation tasks using a ground robot (Turtlebot3 Waffle Pi robot [88, 89]) equipped with a manipulator arm with 4 DOFs (OpenManipulator-X[90]). Unlike Section V-A, the robot is allowed to ask for help, whenever needed, as determined by CP with $\alpha=0.05$. The environment consists of two cans of Coke, one bottle of water, and one pen (see Fig. 1). The robot knows the exact position of each object but the obstacle-free space of the environment is unknown. As a result, the robot is not aware a priori if there is any object that cannot be reached due to blocking obstacles. We use existing navigation and sensing stacks [91] for Turtlebots as well as the MoveIt! [92] toolbox for manipulation control. In the following experiments, we consider GPT 3.5 as the LLM model. The demonstrations are provided in the supplementary video in [86].
Figure 5: Reacting to Infeasible Sub-tasks: As the robot navigates towards Coke $1$, it builds the occupancy grid map of the environment allowing it to realize that Coke $1$ is not accessible. Once this happens, the robot asks for help from the LTL planner.
Reacting to the Unknown Environment: First, we consider the task $\phi=\Diamond(\pi_{1}\vee\pi_{2})$ where $\pi_{1}$ means ‘Deliver Coke $1$ to ${\mathbf{x}}_{E}$’ and $\pi_{2}$ means ‘Deliver Coke $2$ to ${\mathbf{x}}_{E}$ using the environment described as in Fig. 1. This task requires either Coke $1$ or $2$ to be delivered to ${\mathbf{x}}_{E}$. Initially, the LTL planner selects $\pi_{1}$ as $\pi_{\text{next}}$. As the robot navigates the environment to reach Coke $1$, it builds an occupancy grid map of the environment that is used, as in [19], to reason about whether the object is blocked by obstacles or not (see also Fig. 5). Once the robot realizes that Coke $1$ is not reachable, it requests help from the LTL planner. In response to that request, the LTL planner generates an alternative sub-task, modeled by $\pi_{2}$, that is eventually accomplished by the robot. Assistance from a human operator was never requested in this case study since all constructed prediction sets were singletons.
Reacting to Ambiguous Sub-Tasks: Second, we consider a task $\phi=\Diamond\pi_{1}\wedge\Diamond\pi_{2}$ where both $\pi_{1}$ and $\pi_{2}$ mean ‘Bring a drink to location ${\mathbf{x}}_{C}$’. Observe that these APs are ambiguous as both water and Coke qualify as drinks. Once the LTL planner generates the sub-task $\pi_{1}$, the LLM selects the action ‘go to the Coke $2$ location’. However, the prediction set includes two actions ‘go to the bottle-of-water location’ and ‘go to the Coke $2$ location’. Interestingly, we did not specify in the prompt that both water and Coke qualify as drinks. In this case, the robot asks for help from the LTL planner. The LTL planner cannot provide assistance as there are no alternative sub-tasks to make mission progress. Thus, the robot next seeks help from a user. Once the user selects the desired action and $\pi_{1}$ is satisfied, the LTL planner generates the next sub-task $\pi_{2}$ and the above process repeats until the mission is accomplished.
V-D Effect of Predicate Definitions on Planning Performance
In this section, first we show that a fixed mission can be represented by various LTL-NL formulas that differ in the definition of the NL-based predicates. Then, we demonstrate how various definitions of these predicates, resulting in equivalent LTL-NL formulas, can affect performance of HERACLEs.
Specifically, we consider a simple mission requiring to deliver objects A, B, C, and D to desired destinations in any order. We can express this task using an LTL-NL formula with four APs: $\phi_{1}=(\Diamond\pi_{A})\wedge(\Diamond\pi_{B})\wedge\Diamond(\pi_{C})\wedge\Diamond(\pi_{D})$. In $\phi_{1}$, $\pi_{i}$ is true when the object $i\in\{A,B,C,D\}$ is delivered to the intended destination. An equivalent LTL-NL formula is $\phi_{2}=(\Diamond\pi_{A,B})\wedge\Diamond(\pi_{C})\wedge\Diamond(\pi_{D})$ where the only difference is that now we have defined a predicate $\pi_{A,B}$ that is true if both $A$ and $B$ are delivered to their destinations in any order. In a similar way, we can define the equivalent LTL-NL formulas $\phi_{3}=(\Diamond\pi_{A,B,C})\wedge\Diamond(\pi_{D})$ and $\phi_{4}=(\Diamond\pi_{A,B,C,D})$. Observe that formula $\phi_{4}$ is the most ‘user-friendly’ one as it has the fewest predicates and temporal operators. However, this results in increasing the number of sub-tasks that each predicate models. This in turn requires the LLM to reason about more complex task requirements that would otherwise have been handled by the symbolic planner. For instance, in $\phi_{4}$, the symbolic planner generates the subtask $\pi_{A,B,C,D}$ and the LLM needs to design a feasible plan for it. Increasing the complexity of the sub-tasks modeled by the predicates may in turn result in higher help rates. Specifically, we run HERACLEs using the open-source Llama 2-13b model to compute plans for each LTL-NL formula. When $0.91-\alpha=$ 0.9, the average help rates over $10$ trials for $\phi_{1}$, $\phi_{2}$, $\phi_{3}$, and $\phi_{4}$ were $10\%$, $35\%$, $50\%$, and $85\%$ respectively; each trial differs in the calibration dataset. Notice that in case of $\phi_{4}$, HERACLEs is essentially equivalent to [41] since the symbolic planner cannot perform any task decomposition by construction of $\phi_{4}$.
V-E Comparisons of LTL-NL and LTL Formulas
In this section, we provide examples to demonstrate that LTL-NL formulas are more user-friendly than their corresponding LTL formulas defined over the system states ${\mathbf{p}}(t)$ capturing the same mission requirements. As a performance metric, we employ the ‘length’ of the specification, i.e., the total number of atomic predicates and temporal and logical operators. In what follows, we select an LTL-NL formula for each case study considered in Section V-A and we manually convert it into the respective LTL formula. First, consider the LTL-NL formula $\phi=\Diamond(\pi_{1})$ from Case Study I, where the NL-based predicate $\pi_{1}$ is true if the robot delivers a bottle of water to location $A$. This formula only has $1$ temporal/logical operator and $1$ predicate. Using system-based predicates, the same task will be written as $\phi=\Diamond(\pi_{1}^{\prime}\wedge(\Diamond\pi_{2}^{\prime}\wedge(\Diamond\pi_{3}^{\prime}\wedge(\Diamond\pi_{4}^{\prime}))))$ where $\pi_{1}^{\prime}$ is true if the robot position is close enough to the bottle of water, $\pi_{2}^{\prime}$ is true if the robot grabs the bottle successfully, $\pi_{3}^{\prime}$ is true if the robot position is close enough to location $A$, and $\pi_{4}^{\prime}$ is true if the robot puts down the bottle of water. Second, consider consider the LTL-NL formula $\phi=\Diamond(\pi_{1})\wedge\Diamond(\pi_{2})$ from Case Study II where both $\pi_{1}$ and $\pi_{2}$ model delivery tasks as before. This formula has $3$ logical/temporal operator and $2$ predicates. Expressing the same task using system-based predicates as before would result in $15$ logical/temporal operators and $8$ predicates, i.e., $\phi=\Diamond(\pi_{1}^{\prime}\wedge(\Diamond\pi_{2}^{\prime}\wedge(\Diamond\pi_{3}^{\prime}\wedge(\Diamond\pi_{4}^{\prime}))))\wedge\Diamond(\pi_{5}^{\prime}\wedge(\Diamond\pi_{6}^{\prime}\wedge(\Diamond\pi_{7}^{\prime}\wedge(\Diamond\pi_{8}^{\prime}))))$. Third, we select the following formula from Case Study III: $\phi=\Diamond\pi_{1}\wedge\Diamond\pi_{2}\wedge\Diamond\pi_{3}\wedge\Diamond\pi_{4}\wedge(\neg\pi_{4}{\mathcal{U}}\pi_{1})$. This LTL-NL formula has $10$ logical/temporal operators and $4$ predicates, where the corresponding LTL formula over system-based predicates requires $46$ logical/temporal and $16$ predicates. Observe that the difference in the length of LTL and LTL-NL formulas becomes more pronounced as the complexity of the mission requirements increases. Finally, another reason why LTL-NL formulas are more user-friendly than their corresponding LTL formulas is that their predicates can be defined using natural language instead of low-level system configurations.
VI Conclusions, Limitations, and Future Work
Summary: In this paper, we proposed HERACLEs, a new neuro-symbolic planner for LTL-NL tasks. We showed, both theoretically and empirically, that it can achieve desired mission success rates due to a conformal interface between the symbolic planner and the LLM. We also provided comparative experiments showing that it outperforms LLM-based planners in terms of planning accuracy and help rates as well as examples to demonstrate its user-friendliness compared to symbolic planners.
Limitations: Key limitations of our work are the assumptions of (i) perfect execution of robot skills; (ii) faithful users providing correct decisions; and (iii) generating i.i.d. validation and calibration data.
Future Work: Our future work will focus on addressing these limitations. Specifically, (i) can be addressed as discussed in Remark IV.3 while (ii) can be addressed by integrating human errors in CP. To address the limitation in (iii), we will leverage robust CP frameworks [85]. Our future work will also focus on designing translation frameworks that can convert NL instructions into the desired LTL-NL specifications with translation correctness guarantees.